OpenStudy (mlparks13):

Solve for x: 4|2x − 2| + 10 = 34. x = −4, x = 2 x = 4, x = −4 x = 2, x = −2 x = 4, x = −2 Please answer i will reward right answer with medal

OpenStudy (danjs):

isolate the absolute value quantity on one side by itself first

OpenStudy (mlparks13):

ok

OpenStudy (mlparks13):

put 4|2x − 2| by it self

OpenStudy (mlparks13):

put 4|2x − 2| by it self

OpenStudy (danjs):

just the absolute value part.

OpenStudy (danjs):

4|2x − 2| + 10 = 34 |2x − 2| = 6

OpenStudy (danjs):

remember the graph of absolute value of x, the v shape...  notice for each value, it can intersect the graph 2 times and have two answers, y = abs(x), y=x or y=-x